# Autonomous Behavior and Whole-Brain Dynamics Emerge in Embodied Zebrafish Agents with Model-based Intrinsic Motivation

### v1

@@ Abstract
Autonomy is a hallmark of animal intelligence, enabling adaptive and intelligent behavior in complex environments without relying on external reward or task structure. Existing reinforcement learning approaches to exploration in sparse reward and reward-free environments, including class of methods known as intrinsic motivation, exhibit inconsistent exploration patterns and thus fail to produce robust autonomous behaviors observed in animals. Moreover, systems neuroscience has largely overlooked the neural basis of autonomy, focusing instead on experimental paradigms where animals are motivated by external reward rather than engaging in unconstrained, naturalistic and task-independent behavior. To bridge these gaps, we introduce a novel model-based intrinsic drive explicitly designed to capture robust autonomous exploration observed in animals. Our method (3M-Progress) motivates naturalistic behavior by tracking divergence between the agent’s current world model and an ethological prior. We demonstrate that artificial embodied agents trained with 3M-Progress capture the explainable variance in behavioral patterns and whole-brain neural-glial dynamics recorded from autonomously-behaving larval zebrafish, introducing the first goal-driven, population-level model of neural-glial computation. Our findings establish a computational framework connecting model-based intrinsic motivation to naturalistic behavior, providing a foundation for building artificial agents with animal-like autonomy.
@@ 1 Introduction
Animals exhibit remarkable autonomy, navigating complex environments through self-directed, internally driven behaviors rather than solely responding to external rewards or immediate physiological needs. Unlike typical artificial agents designed to optimize explicit, predefined task objectives in well-defined problem settings, animals intrinsically explore and adapt in open-ended, naturalistic environments where goals are neither clear nor stable. This capacity for autonomous behavior allows organisms, including humans, to flexibly engage in abstract thinking, exploratory learning, and innovative problem-solving. Central to this autonomy is the ability to generate intrinsic goals, seek novel experiences, and continuously refine internal models of the world that inform future actions. Understanding the interplay between extrinsic and intrinsic motivations remains a fundamental challenge in both systems neuroscience and artificial intelligence, particularly for building robust agents capable of sustained autonomy.
@@
Figure 1: Simulation of zebrafish embodiment in varied virtual environments. A) The base 6-link swimmer [1] in a virtual environment with passive fluid force. B) The agent controls the torque exerted by motors at each joint (5 total degrees of freedom) to swim and navigate its environment. C) A custom cosmetic skin to mimic the appearance of larval zebrafish. D) Custom virtual environment to match the experimental parameters of the open loop protocol Mu et al. [2]. The root joint located at the head is fixed during training.
@@
However, current intrinsic motivation methods for exploration in reinforcement learning (RL) —such as curiosity-driven exploration [3, 4, 5, 6]—often exhibit fragile and inconsistent behavioral patterns due to difficulties distinguishing between controllable and uncontrollable stimuli, coping with non-stationary environments, and avoiding the “white noise problem” [7]. For example, agents driven by prediction-error alone gets locked into exploring inherently unpredictable or irrelevant aspects of the environment, hindering their ability to develop robust exploratory behaviors [3]. Moreover, existing methods typically fail to produce structured and stable behavioral transitions that characterize genuine autonomous behaviors observed in animals and children [8].
@@
At the same time, systems neuroscience has historically overlooked the cellular basis of autonomy, favoring experimental paradigms centered on explicit external rewards to motivate task-dependent behaviors rather than free, unconstrained exploration. As a result, the cellular mechanisms underlying naturalistic autonomy—particularly those involving whole-brain interactions—remain poorly understood. However, growing experimental, computational, and theoretical evidence suggests that non-neuronal cells, especially astrocytes, play a critical role in the generation of intelligent behavior [2, 9, 10, 11, 12, 13, 14]. Owing to their close association with neurons, distinctive connectivity (with a single astrocyte capable of interfacing with up to a million nearby synapses), and ability to perform integrative computations using a hierarchy of timescales, astrocytes are well-positioned to support adaptive, naturalistic, goal-directed behavior [15, 16, 17, 18, 19, 20].
@@
To investigate this question, we study autonomy in larval zebrafish—a uniquely valuable animal model due to their optical transparency, which affords whole-brain calcium imaging via light-sheet microscopy [21]. Given that the cellular basis of autonomous behavior remains largely unknown, whole-brain imaging allows us to search across the entire recorded population of approximately 300,000 cells (including 130,000 neurons and 130,000 astrocytes) in aligning our model with brain activity [2]  [^1] The dataset we use is open-source and freely available here.. We examine a particular cognitive behavior known as futility-induced passivity, observed across the entire animal spectrum [22, 23, 2, 24]. In zebrafish, futility-induced passivity occurs when their active swimming efforts in virtual arenas fail to simulate optic flow. Mu et al. [2] offers rigorous experimental evidence suggesting this transition to passivity is driven neuron-glial interactions in the lateral medulla oblongata (L-MO) that detect and accumulate sensorimotor feedback error. Noradrenergic (NE) neurons signal swim failures which slowly activate radial astrocyte with extending processes to the NE cluster of L-MO. At a critical level of intracellular calcium, astrocyte processes downstream of L-MO activate GABAergic neurons that suppress swimming. By extensive glia manipulation experiments, Mu et al. [2] demonstrate astrocytes are causally required for reliable futility-induced passivity across animal subjects, underscoring the essential role of astrocytes in neural computation and behavioral state modulation.
@@
In this work, we address the challenge of linking neural-glial computations to the robust autonomous behavior observed in embodied settings by introducing a novel intrinsic reinforcement learning method termed Model-Memory-Mismatch Progress (3M-Progress). Our approach leverages an internal model that continually compares an agent’s online memory formed by current sensory experience against an ethologically relevant prior memory; intrinsic reward then reinforces transitions between behaviors that where mismatch between memories is maximized relative to its temporal history. 3M-Progress is uniquely capable of producing stable, ethologically-relevant behavioral transitions among several standard intrinsic motivation methods methods in reinforcement learning.
@@
By training embodied agents with 3M-Progress, we successfully replicate both the robust behavioral transitions observed in autonomously-behaving zebrafish and whole-brain neural-glial dynamics associated with futility-induced passivity. Our agent captures nearly all of the variance in neural and astrocytic activity, marking the first predictive and normative computational model of neural-glial interactions driven by intrinsic motivation.
@@ 2 Related Work
Although glial cells—especially astrocytes—are increasingly recognized as crucial to adaptive brain function [2, 9, 17, 14, 19, 20, 11, 25, 26], computational models of neural-glial interactions remain underdeveloped [27]. Existing models of neuron-astrocyte dynamics typically fall into two categories: phenomenological models that reproduce specific experimental findings like calcium oscillations or epileptic activity [28, 29, 30, 31], and simplified, “bottom-up” mathematical models that explore theoretical principles based on astrocytes’ unique morphology and anatomy [32, 33, 34, 35, 12, 13]. While important, these models are not directly applicable to our setting because they are (a) not yet directly trained on ethological tasks, embodied, or quantitatively validated against real brain data, and (b) typically focused on a single astrocyte, rather than a population of astrocytes. In contrast, we adopt a “top-down” approach: we train a general-purpose recurrent architecture to control an embodied agent to perform ethologically-relevant tasks and find that this imposes strong constraints on the learned representatsion, allowing us to identify units whose activity patterns closely match those observed in the neurons and astrocytes of larval zebrafish.
@@
Existing methods for curiosity-driven exploration have primarily been evaluated in constrained environments with either handcrafted object-centric inputs [6], or externally-defined rewards such as Atari games (e.g., Random Network Distillation (RND) [4], Intrinsic Curiosity Module (ICM) [3], or Disagreement-based methods [5]) or non-embodied settings like reinforcement learning from human feedback (RLHF) with language agents [36]. While some recent works, such as LEXA [37] or Dreamer [38, 39], extend intrinsic curiosity to robotic control tasks using visual inputs, these methods still focus on exploration relative to extrinsic task-defined rewards. In contrast, our study investigates completely open-ended autonomous behavior, explicitly lacking any extrinsic rewards. Furthermore, unlike prior intrinsic motivation methods that often operate on low-dimensional or handcrafted state representations, our approach is trained end-to-end directly from high-dimensional, biologically realistic sensory streams, providing an ethologically relevant exploration setting more akin to that experienced by biological organisms.
@@
In the past few years, several works have recently leveraged embodied AI to bridge computational models with neuroscience. Examples include virtual animal models such as Merel et al.’s virtual rodent [40, 41], which facilitates grounded studies of motor control by replicating rodent motor behaviors across various tasks; Janelia’s biomechanically-detailed virtual fly, capable of diverse realistic locomotion driven by reinforcement learning controllers [42]; and Zador et al.’s Embodied Turing Test position paper [43], which emphasizes developing AI models whose sensorimotor capabilities rival those of their biological counterparts.
Our work extends these directions by providing the first predictive and normative computational formalization of neural-glial interactions within embodied agents, trained explicitly via intrinsically-driven exploration. By leveraging biologically-realistic, open-ended scenarios rather than simplified tasks, our approach yields a naturalistic theory of neural-glial computation, offering insights into evolutionary pressures underlying the detailed heterogeneity observed at level of individual neural and glial cells.
@@ 3 Methods
Animals are physically coupled to the environment through their embodiment; this coupling is often referred to as the sensory-motor or perception-action feedback loop. Physical embodiment imposes strong constraints on both the sensory stream from which an agent learns meaningful representations of the world and the actuation system by which the agent manifests behavior. Following this top-down view of biological systems, we constructed an embodied agent and custom virtual environment in the MuJoCo physics engine [44], specifically designed after the ethology of the zebrafish (Figure 1). Leveraging the procedurally generated $n$-link swimmer and built-in inertial fluid model from the Deepmind Control Suite (dm-control)  [1], we construct an ethological environment (Figure 1A-C) in which the agent can freely behave in the presence of both passive and active fluid currents, similar to the gentle water sources to which zebrafish are native. To evaluate our agent in the futility-induced passivity task, we construct a second environment closely matching the open-loop experimental protocol in Mu et al. [2] (Figure 1D). In this environment, agents passively experience a high-contrast grating moving away from the egocentric point-of-view, which simulates backward motion. In the closed-loop condition, the agent is head-free and can learn a positional-homeostasis policy to counteract the perceived backward flow. In the open-loop condition, the agent is head-fixed and its swim commands produce no movement. The passive speed of the moving grating, its colors, and sizing relaitve to the zebrafish body were determined from experimental parameters in Mu et al. [2]. These environments provide sensory and actuator configurations that closely match the basic structure of free swimming as well as the head-fixed experimental protocol both during model training and model-brain evaluation, which facilitates realistic and meaningful comparison between artificial and biological agents.
Zebrafish Agent Architecture and Policy Training
@@
The zebrafish agent is equipped with a recurrent sensory-cognitive architecture to support perception and action in complex environments. Because the autonomous zebrafish behavior recorded by Mu et al. [2] is primarily driven by visual input, we restrict the sensory stream to a vision encoder operating on images with similar resolution to the visual acuity of zebrafish [45]. Sensory features are used by three distinct cognitive networks with modularized objectives. The core and policy modules are implemented as an actor-critic architecture using Long Short-term Memory (LSTM) networks [46] followed by feedforward decoders trained end-to-end with Proximal Policy Optimization (PPO) [47]. Although PPO is a model-free reinforcement learning algorithm, the intrinsic drive module (IDM) learns a world model through online experience that approximates the world state-action transition dynamics in sensory feature-space (as described below). However, this internal model functions only to generate intrinsic reward and is not used for planning or training the policy.
@@
Figure 2: Agent architecture and 3M-Progress. A) Egocentric visual input ($I_{t})$ is ecnoded via a small residual network. Proprioceptive state observations ($J_{t}$) are encoded via a small multi-layer perceptron with shortcut paths to both the core and policy module. Sensory features are passed into recurrent LSTM core ($h_{t}^{c}$) and policy ($h_{t}^{\pi}$) modules that learn a state value function and stochastic policy, respectively. The intrinsic drive module consists of a small multi-layer perceptron that parameterizes a forward dynamics model on sensory features observed from an environment with dynamics $T_{\mathrm{world}}$. B) 3M-Progress uses two memories created from environments with differing transition dynamics. Divergence between the ethological prior $\omega_{\theta}$ and the current world-model $\omega_{\theta^{\prime}}$ defines 3M, which is then is used as input to leaky integrator $\hat{\epsilon}$ to generate intrinsic reward $r_{t}^{i}$.

### v2

# Intrinsic Goals for Autonomous Agents: Model-Based Exploration in Virtual Zebrafish Predicts Ethological Behavior and Whole-Brain Dynamics

@@ Abstract
Autonomy is a hallmark of animal intelligence, enabling adaptive and intelligent behavior in complex environments without relying on external reward or task structure. Existing reinforcement learning approaches to exploration in reward-free environments, including a class of methods known as model-based intrinsic motivation, exhibit inconsistent exploration patterns and do not converge to an exploratory policy, thus failing to capture robust autonomous behaviors observed in animals. Moreover, systems neuroscience has largely overlooked the neural basis of autonomy, focusing instead on experimental paradigms where animals are motivated by external reward rather than engaging in ethological, naturalistic and task-independent behavior. To bridge these gaps, we introduce a novel model-based intrinsic drive explicitly designed after the principles of autonomous exploration in animals. Our method (3M-Progress) achieves animal-like exploration by tracking divergence between an online world model and a fixed prior learned from an ecological niche. To the best of our knowledge, we introduce the first autonomous embodied agent that predicts brain data entirely from self-supervised optimization of an intrinsic goal—without any behavioral or neural training data—demonstrating that 3M-Progress agents capture the explainable variance in behavioral patterns and whole-brain neural-glial dynamics recorded from autonomously behaving larval zebrafish, thereby providing the first goal-driven, population-level model of neural-glial computation. Our findings establish a computational framework connecting model-based intrinsic motivation to naturalistic behavior, providing a foundation for building artificial agents with animal-like autonomy.
@@ 1 Introduction
Animals exhibit remarkable autonomy, navigating complex environments through self-directed, internally driven behaviors rather than solely responding to external rewards or immediate physiological needs. Unlike typical artificial agents designed to optimize explicit, predefined task objectives in well-defined problem settings, animals intrinsically explore and adapt in open-ended, naturalistic environments where goals are neither clear nor stable. This capacity for autonomous behavior allows organisms, including humans, to flexibly engage in abstract thinking, exploratory learning, and innovative problem-solving. Central to this autonomy is the ability to generate intrinsic goals, seek novel experiences, and continuously refine internal models of the world that inform future actions. Understanding the interplay between extrinsic and intrinsic motivations remains a fundamental challenge in both systems neuroscience and artificial intelligence, particularly for building robust agents capable of lifelong autonomy.
@@
Figure 1: Simulation of the zebrafish agent in a physics-based virtual environment. A) The 6-link embodiment geometry [1] in a environment with dynamic fluid forces. B) The agent controls the torque exerted by motors at each joint (5 DoF) to swim and navigate its environment. C) A custom cosmetic skin to mimic the appearance of larval zebrafish. D) A virtual environment matching the experimental parameters of the open loop protocol Mu et al. [2]. The root joint located at the head is fixed during training.
@@
However, existing methods in reinforcement learning for intrinsic motivation using world-models —such as curiosity-driven exploration and variants [3, 4, 5, 6]—exhibit inconsistent behavioral patterns due to difficulties distinguishing between controllable and uncontrollable stimuli, coping with non-stationary environments, and avoiding the “noisy TV” problem [7]. For example, agents driven by reinforcing prediction-error alone gets locked into pursuing inherently unpredictable or irrelevant aspects of the environment, hindering their ability to develop robust exploratory behaviors [3]. Moreover, existing methods typically fail to produce structured and stable behavioral transitions that characterize genuine autonomous behaviors observed in animals and children [8].
@@
At the same time, systems neuroscience has historically overlooked the cellular basis of autonomy, favoring experimental paradigms centered on explicit external rewards to motivate task-dependent behaviors rather than free, unconstrained exploration. As a result, the cellular mechanisms underlying naturalistic autonomy—particularly those involving whole-brain interactions—remain poorly understood. Growing experimental, computational, and theoretical evidence suggests that non-neuronal cells, especially astrocytes, play a critical role in the generation of intelligent behavior [2, 9, 10, 11, 12, 13, 14]. Owing to their close association with neurons, distinctive connectivity (with a single astrocyte capable of interfacing with up to a million nearby synapses) and ability to perform integrative computations using a hierarchy of timescales, astrocytes are well-positioned to support adaptive, naturalistic, goal-directed behavior [15, 16, 17, 18, 19, 20].
A Dataset for Animal Autonomy
@@
To investigate these questions, we study autonomy in larval zebrafish—a uniquely valuable animal model due to their optical transparency, which affords whole-brain calcium imaging via light-sheet microscopy [21]. Given that the cellular basis of autonomous behavior remains largely unknown, whole-brain imaging allows us to search across the entire recorded population of over 250,000 cells (roughly 125K neurons and 125K astrocytes) to measure how our proposed model aligns with brain activity [2]  [^1] The dataset we use is open-source and freely available here.. We examine a particular cognitive behavior known as futility-induced passivity, an experimentally induced state of "giving up" observed across the animal spectrum [2, 22, 23, 24]. In zebrafish, futility-induced passivity occurs when their swimming effort in a virtual environment fails to simulate optic flow, a key visual feedback signal associated with motor-action under ethological circumstances. After some time in the passive state, zebrafish reattempt to swim–an exploratory behavior to assess whether the open-loop dynamics of the environment have changed. This cycle between active and passive states continues over the course of a trial.
Mu et al. [2] offers rigorous experimental evidence suggesting this transition to passivity is driven neuron-glial interactions in the lateral medulla oblongata (L-MO) that detect and accumulate sensorimotor feedback error. Noradrenergic (NE) neurons signal swim failures which slowly activate radial astrocyte with extending processes to the NE cluster of L-MO. At a critical level of intracellular calcium, astrocyte processes downstream of L-MO activate GABAergic neurons that suppress premotor neurons relevant for swimming. With extensive manipulation experiments, Mu et al. [2] demonstrate that futility-induced passivity is not due to fatigue, struggle, or positional homeostasis. Specifically, when visual feedback from closed-loop trials was replayed during the open-loop condition, zebrafish still transitioned to a passive state. This indicates the behavior is mediated by an internal model that is used to detect active sensorimotor feedback error, providing strong evidence for model-based intrinsic motivation.
Contributions
@@
In this work, we study neural-glial computations and their relationship to autonomous animal behavior by training an embodied zebrafish agent with model-based intrinsic motivation and studying its emergent behavior and internal representations. We introduce a novel intrinsic reinforcement learning algorithm termed Model-Memory-Mismatch Progress (3M-Progress) alongside a virtual environment that captures the fundamental physics of the zebrafish embodiment. Our approach leverages an internal model that continually compares the agent’s online memory formed by its current sensory experience against an ethologically relevant prior memory; intrinsic reward then reinforces transitions between behaviors that maximize the divergence between memories relative to its temporal history. 3M-Progress is uniquely capable of producing stable, ethologically-relevant behavioral transitions among several state-of-the-art exploration algorithms in reinforcement learning.
@@
By training embodied agents with 3M-Progress, we successfully replicate both the robust behavioral patterns and whole-brain neural-glial dynamics in autonomously-behaving zebrafish. Capturing nearly all of the variance in neural and astrocytic activity, this marks the first predictive and normative model of neural-glial computation. Our agent was not trained on any behavioral or neural data, and thus represents the first autonomous embodied agent that predicts brain data completely from optimizing a self-supervised, intrinsic goal. To summarize, our technical contributions are:
  • 3M-Progress, a novel intrinsic reward that leverages an ecological dynamics prior to guide exploration in new environments.
  • Emergence of an autonomous behavior known as futility-induced passivity in 3M-Progress agents, closely matching larval zebrafish behavior.
  • Alignment between whole-brain calcium response and 3M-Progress agents, providing the first goal-driven model of neural-glial computation.
  • A general modeling perspective positioning intrinsic reinforcement learning as a computational framework for understanding autonomy in animals.
@@ 2 Related Work
Although glial cells—especially astrocytes—are increasingly recognized as crucial to adaptive brain function [2, 9, 17, 14, 19, 20, 11, 25, 26], computational models of neural-glial interactions remain underdeveloped [27]. Existing models of neuron-astrocyte dynamics typically fall into two categories: phenomenological models that reproduce specific experimental findings like calcium oscillations or epileptic activity [28, 29, 30, 31], and simplified, “bottom-up” mathematical models that explore theoretical principles based on astrocytes’ unique morphology and anatomy [32, 33, 34, 35, 12, 13]. While important, these models are not directly applicable to our setting because they are (a) not yet directly trained on ethological tasks, embodied, or quantitatively validated against real brain data, and (b) typically focused on a single astrocyte rather than a population of astrocytes. In contrast, we adopt a “top-down” approach: we train a general-purpose recurrent architecture to control an embodied agent to perform ethologically-relevant behavior and find that this imposes strong constraints on the learned representations, allowing us to identify units whose activity patterns closely match those observed in neurons and astrocytes of larval zebrafish.
@@
Exploration using self-supervised world-models has demonstrated promising success in several standard reinforcement learning domains, and even more recently in language modeling  [36]. Methods like learning progress  [6] and Random Network Distillation (RND)  [4] were primarily evaluated using either handcrafted object-centric or state observations with low-dimensionality embodiments, limiting the applicability to pixel-based environments or continuous control. The Intrinsic Curiosity Module (ICM)  [3] uses a pixel-encoder trained with an inverse-dynamics loss to predict features rather than raw states, but was evaluated on discrete pixel-based environments like Doom and Atari. While some recent works, such as LEXA [37] or Plan2Explore [38], extend intrinsic curiosity (specifically, Disagreement  [5]) to continuous control from visual inputs, success of the exploration policy is defined relative to downstream task generalization. Since these methods do not independently evaluate the quality of the exploration policy, it remains unclear whether these algorithms are powerful enough to learn complex behaviors. In contrast, we investigate completely open-ended autonomous behavior in reward-free, continuous MDPs with high-dimensional observations, turning the focus on the ability of the exploration algorithm to develop ethological, interpretable behaviors independent of any downstream task.
@@
Several works have leveraged embodied AI to bridge computational models with neuroscience, including virtual animal models such as the virtual rodent [39, 40], which facilitates grounded studies of motor control by replicating rodent motor behaviors across various tasks using imitation learning; the virtual fruit fly, a biomechanically detailed model matching both the visual system and basic flight capacity used to study a diverse range of behaviors driven by imitation learning [41, 42, 43]; the OpenWorm project [44], a biophysically accurate simulation of the C-elegan nematode, but has yet to be combined with deep learning and task-optimization; and Zador et al.’s Embodied Turing Test position paper [45], which emphasizes developing AI models whose sensorimotor capabilities rival those of their biological counterparts. There are several existing works that apply task-optimization to control details musculoskeletal models [46, 47, 48], use robots to implement and validate neural circuits in zebrafish [49], and model animal-like social behavior or object perception using digital twins [50, 51].
Our work extends these directions by providing the first predictive and normative computational formalization of neural-glial interactions in embodied agents, thereby validating a circuit model recently proposed by Mu et al. [2]. Most importantly, our model is trained entirely via intrinsically-motivated exploration, unlike previous approaches that constrain behavior and their resulting neural representations by supervised learning  [39, 40, 41]. To the best of our knowledge, this marks the first completely autonomous, embodied agent model of behavioral and brain data in neuroscience, pointing towards a promising computational framework for understanding naturalistic, task-independent behavior in biological systems using intrinsic reinforcement learning.
@@ 3 Methods
Animals are physically coupled to the environment through their embodiment; this coupling is often referred to as the sensory-motor or perception-action feedback loop. Physical embodiment imposes strong constraints on both the sensory stream from which an agent learns meaningful representations of the world and the actuation system by which the agent manifests behavior. Following this top-down view of biological systems, we construct an embodied agent and custom virtual environment in the MuJoCo physics engine [52] specifically designed after the ethology of the zebrafish (Figure 1). Leveraging the procedurally generated $n$-link swimmer and built-in inertial fluid model from the Deepmind Control Suite (dm-control)  [1], we construct an ethological environment (Figure 1A-C) in which the agent can freely behave in the presence of both passive and active fluid currents, similar to the dynamic water environments to which zebrafish are native. To evaluate our agent in the futility-induced passivity task, we construct a second environment closely matching the open-loop experimental protocol in Mu et al. [2] (Figure 1D). In this environment, agents passively experience a high-contrast grating moving away from the egocentric point-of-view, which simulates backward motion. In the closed-loop condition, the agent is head-free and can learn a positional-homeostasis policy to counteract the perceived backward flow. In the open-loop condition, the agent is head-fixed and its swim commands produce no movement. The passive speed of the moving grating, its colors, and sizing relative to the zebrafish body were determined from experimental parameters in Mu et al. [2]. These environments provide sensory and actuator configurations that closely match the basic structure of free swimming as well as the head-fixed protocol, facilitating a meaningful comparison between artificial and biological agents.
Agent Design
@@
The zebrafish agent is equipped with a recurrent sensory-cognitive architecture to support perception and action in continuous, high-dimensional environments. Because the autonomous zebrafish behavior recorded by Mu et al. [2] is primarily driven by visual input, we restrict the sensory stream to a vision encoder operating on images with similar resolution to the visual acuity of zebrafish [53]. Sensory features are used by three distinct cognitive networks with modularized objectives. The core and policy modules are implemented as an actor-critic architecture using Long Short-term Memory (LSTM) networks [54] followed by feedforward decoders trained end-to-end with Proximal Policy Optimization (PPO) to output torques that are used as the control input to the agent’s motors [55]. Specific agent implementation details can be found in Appendix E. Although PPO is a model-free reinforcement learning algorithm, the intrinsic drive module (IDM) learns a world model through online experience that approximates the world state-action transition dynamics in sensory feature-space (as described below). However, this internal model functions only to generate intrinsic reward and is not used for planning or model-based control.
@@
Figure 2: Agent architecture and 3M-Progress. A) Egocentric visual input ($I_{t})$ is encoded via a small residual network $\phi_{I}$. Proprioceptive state observations ($J_{t}$) are encoded via a small multi-layer perceptron $\phi_{J}$ with shortcut paths to both the core and policy module. Sensory features are passed into recurrent LSTM core ($h_{t}^{c}$) and policy ($h_{t}^{\pi}$) modules that learn a state value function and stochastic policy, respectively. The intrinsic drive module consists of a small multi-layer perceptron that parameterizes a forward dynamics model on sensory features observed from an environment with dynamics $T_{\mathrm{world}}$. B) 3M-Progress uses two memories created from environments with differing transition dynamics. Divergence between the ethological prior $\omega_{\theta}$ and the current world-model $\omega_{\theta^{\prime}}$ defines 3M, which is then is used as input to leaky integrator $\hat{\epsilon}$ to generate intrinsic reward $r_{t}^{i}$.